1.state the general solution to the equation sec2x=-3
2. use an algebraic approach to solve the equation cos3x=-1/2

Answer by stanbon(75887)

state the general solution to the equation sec2x=-3
-----
If sec(2x) = -3
then cos(2x) = -1/3
Find the arccos(-1/3) = 109.47 + 720n degrees OR -109.47 + 720n degrees
2x = 109+720n OR 2x = -109.47+720n
x = 54.5+360n degrees OR x = -54.4+360n degrees
----------------------------
2. use an algebraic approach to solve the equation cos3x=-1/2
3x = arccos(-1/2)
3x = 120 degrees OR 240 degrees
x = 40 degrees OR 80 degrees
======================
Cheers,
Stan H.